OpenStudy (anonymous):

The ratio of the radius of circle A to the radius of circle B is 3:5. The ratio of the radius to circle A to the radius of circle C is 3:4. What is the ratio of the area of circle B to circle C? 5:4 9:16 9:25 15:11 25:16

Parth (parthkohli):

Hints: 1) Radius of circle A = 3x 2) Radius of circle B = 5x 3) Radius of circle C = 4x

Parth (parthkohli):

All you should do is to find\[5x \over 4x \]

OpenStudy (ghazi):

R (A/B)=3/5 and R(A/C)=3/4 therefore R(B/C)=(3/4)/(3/5)= 3/4*5/3=5/4

OpenStudy (ghazi):

ratios of area= ratios of square of radius

Parth (parthkohli):

Obviously, the radius of circle C won't be 0... so you can cancel.

OpenStudy (ghazi):

(5/4)^2=25/16

Parth (parthkohli):

\[5x \over 4x \]\[5 \over 4 \]

Parth (parthkohli):

Now you have to find the ratios of their squares.

OpenStudy (anonymous):

so the answer would be 5:4?

Parth (parthkohli):

No.

OpenStudy (ghazi):

nope square of that

Parth (parthkohli):

\[{5^2 \over 4^2} \]

OpenStudy (anonymous):

25:16

Parth (parthkohli):

yes